OpenStudy (anonymous):

Rationalize the denominator and simplify. Could someone explain? 5√3 + 2 ------- √3 + 1

OpenStudy (anonymous):

The dotted line means divide

OpenStudy (zela101):

Are you familiar with rationalizing the denominator?

OpenStudy (alexandervonhumboldt2):

this is so easy: multiply by a conjugate: conjugate of sqrt3+1 is sqrt3-1 \[\frac{ 5\sqrt3+2 }{ \sqrt3+1 }=\frac{ 5\sqrt3+2 }{ \sqrt3+1 }*\frac{ \sqrt3-1 }{ \sqrt3-1 }\]

OpenStudy (alexandervonhumboldt2):

\[=\frac{ (5\sqrt3+2)*(\sqrt3-1) }{ 3-1 }=\frac{ 5*3-5+2\sqrt3-2 }{ 2 }=?\]

OpenStudy (alexandervonhumboldt2):

woppers

OpenStudy (alexandervonhumboldt2):

mistake

OpenStudy (alexandervonhumboldt2):

\[\frac{ 5*3-5\sqrt3+2\sqrt3-2 }{ 2 }=?\]

OpenStudy (anonymous):

So I would have to foil it ?

OpenStudy (anonymous):

I know the denominator ends up being 2 because the √3 * √3 = 3 and then subtract 1 = 2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah foil it and simplify

OpenStudy (anonymous):

what do I do with the middle numbers? multiply?

OpenStudy (anonymous):

in the numerator?

OpenStudy (anonymous):

15 - 7√3 - 2 = 13 - 7√2?

OpenStudy (alexandervonhumboldt2):

nope -5sqrt3+2sqrt3=-3sqrt3 thus we have that 13-3sqrt3

OpenStudy (anonymous):

so it would be 13- 3√3/2

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

Ah okay! Thank you so much!